Circular No. 6009

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET BORRELLY (1994l) This comet has been recovered independently by A. C. Gilmore and

P. M. Kilmartin and by G. J. Garradd, as shown below; these positions are precisely consistent with the prediction on MPC 18259 (ephemeris on MPC 23144):

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
June 12.72895 1^{h}47^{m}26.^{s}62 −18^{o}52′31′′.9 16.7 Gilmore
12.75175 1 47 29.71 −18 52 20.3  _{”}
16.72160 1 56 34.33 −18 18 14.7 16.8  _{”}
16.73462 1 56 36.13 −18 18 07.7  _{”}
17.82322 1 59 06.22 −18 08 48.8  McNaught
17.82597 1 59 06.57 −18 08 47.5  _{”}
17.82843 1 59 06.93 −18 08 45.9  _{”}

A. C. Gilmore (Mt. John). Measurer P. M. Kilmartin. R. H. McNaught, G. J. Garradd, D. J. Asher and D. I. Steel (Siding Spring).

SUPERNOVA 1994I IN NGC 5194 J. E. H. Wink, Institut de Radio Astronomie Millimétrique (IRAM),

reports: “SN 1994I was observed at 88 GHz during Apr. 13.83–14.28 UT with the IRAM interferometer. The source was unresolved with a 3′′ beam at α = 13^{h}27^{m}47^{s}.7, δ = +47^{o}26′58′′ (error 0′′.5). The flux was 2.83 mJy (1σ error 0.44 mJy); this was calibrated against the source 1418+546, which measured 0.880 Jy. This flux density, when compared with 15-GHz observations of 15 mJy made at the same time by Pooley and Green (K. Weiler, priv. comm.), indicates a nonthermal spectral index of −0.9 (error 0.1). The 230-GHz detections of 8 mJy reported by Chini et al. (IAUC 5969) are likely contaminated by dust emission from NGC 5194.”

AG DRACONIS The following visual magnitude estimates indicate that this symbiotic

variable has entered a bright outburst: May 9.97 UT, 9.9 (B. H. Granslo, Fjellhamar, Norway); 26.99, 9.8 (Granslo); June 2.99, 9.3 (Granslo); 4.99, 9.3 (G. Poyner, Birmingham, England); 14.63, 8.4 (Y. Takenaka, Kyoto, Japan); 19.094, 8.5 (P. Schmeer, Bischmisheim, Germany).

1994 June 21 Daniel W. E. Green